OpenStudy (anonymous):

A bicycle tire has a diameter of twenty-six inches. If a pebble is stuck in the tire's treads, about how far will the pebble travel with each rotation of the tire? forty point eight two inches eighty-one point six four inches five hundred thirty point six six inches two thousand one hundred twenty-two point six four inches Can someone explain to me how to do this?

OpenStudy (anonymous):

circumference of a circle is given by: \[ c = 2\pi r\] where r is radium \[ r = \frac{diameter}{2} = \frac{26}{2} = 13\] therefor, with each rotation, pebble travels \[ c = 2\pi r = 2 \pi (13) = 26\pi \space m\]

OpenStudy (anonymous):

radius*

OpenStudy (anonymous):

Thanks dude.

OpenStudy (anonymous):

so what is the answer